### v3

# On Leakage of Code Generation Evaluation Datasets

Abstract
In this paper, we consider contamination by code generation test sets, in particular in their use in modern large language models. We discuss three possible sources of such contamination and show findings supporting each of them: (i) direct data leakage, (ii) indirect data leakage through the use of synthetic data and (iii) overfitting to evaluation sets during model selection. To address this, we release LBPP: an uncontaminated new benchmark of 161 prompts with their associated Python solutions. LBPP is released at https://huggingface.co/datasets/CohereForAI/lbpp.
1 Introduction
Code generation has emerged as an important skill for large language models to master. Measuring recent progress in code generation has relied on few, critical benchmarks to judge performance between model families and checkpoints. While many recent sophisticated evaluation datasets have been proposed (Jain et al., 2024; Jimenez et al., 2024), the community largely relies on HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) to judge a new model’s code capability. In fact, all major announcements in 2023-2024 claiming advanced code capabilities—from either academic or industry labs—boast one or both of these datasets. Practically, reporting HumanEval and MBPP is mandatory for a model to report competitive code generation.
However, the importance of these benchmarks has led to a conflict between popularity and utility. Obtaining competitive numbers comes with significant scientific and economic reward—made increasingly easy with the proliferation of public replicas and evaluation harnesses featuring these datasets. However, this prevalence has led to data leakage beyond the original evaluation scope. When this evaluation data contaminates model training, the validity of the metrics as a measure of generalization capability becomes unreliable. If a model is trained on data used for out-of-distribution generalization (or selected for its performance on that data), we break an implicit tenet of how model capability should be measured. We argue that understanding the effect of contamination is critical to accurately interpreting scores on these benchmarks.
In this paper, we review the evidence that most contemporary LLMs are contaminated with data sourced from these two benchmarks. We define contamination here as any procedure leaking datasets during model training such that these datasets are no longer unseen at inference. The most obvious method of contamination is the presence inside training data. Section 3.1 reviews evidence that these benchmarks are widespread in training corpora in original and paraphrased forms. Unfortunately, it is not feasible to manually remove all the corresponding examples from the training corpora and the most common automatic decontamination methods have low recall. Section 3.2 proposes that contamination also occurs indirectly through the use of synthetic data—a widespread paradigm used to increase coding capabilities by generating additional code training tokens for pre-training or fine-tuning. Finally, Section 3.3 argues that checkpoint selection may be overly influenced by these datasets, overfitting to these benchmarks over general-purpose code-oriented generalization.
In this paper, we propose a more challenging Python code generation benchmark: Less Basic Python Problems. LBPP is similar in size to HumanEval and MBPP, but designed to be more complex using model in the loop filtering. LBPP is also designed to share no inspiration or sources with existing training and evaluation datasets, presenting a novel generalization challenge to contemporary LLMs. In Section 4, we observe that SOTA models on HumanEval perform up to $43\%$ worse on LBPP. We contribute LBPP as a genuinely held-out test to measure current code generation capability, and potential overfitting to HumanEval and MBPP.
2 Related Work
HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) remain the most reported results on public leaderboards, but similar datasets exist (Hendrycks et al., 2021; Li et al., 2022). They consist of short and mostly simple (not programming competition level) instructions with completions in Python. These datasets have also been translated into more programming languages (Muennighoff et al., 2023; Cassano et al., 2022), as well as versions with additional tests (Liu et al., 2024b).
Jain et al. (2024) proposes a continuously updated set of interview questions to improve dataset challenge by including harder and novel (unseen) prompts. Jimenez et al. (2024) aims for challenging software engineering problems requiring understanding of entire repositories. In a similar vein, RepoQA (Liu et al., 2024a) and Bug In The Code Stack (Lee et al., 2024) focus on understanding long contexts within code tasks. Zhang et al. (2024) proposes using hidden evaluation sets, however, this approach does not allow inspection of failure cases and requires trusting the quality and correctness of an opaque ‘black-box’ evaluation setup.
Recently, Riddell et al. (2024) analyzed data contamination in popular pretraining datasets: reporting that $12.2\%$ of HumanEval samples are present in The Pile (Gao et al., 2020), and $18.9\%$ in The Stack (Kocetkov et al., 2022). While Riddell et al. (2024) reports “we do not find the performance rankings of the models to change with decontaminated results”, we identify the ranking between models to vary between contaminated and uncontaminated evaluation datasets (see Table 2).
3 Possible sources of contamination
We provide three hypotheses, with respective evidence, on why existing evaluation datasets are leaked and models may already be over-optimized towards existing leaked benchmarks.
3.1 Direct data leakage
The most obvious reason is the simplest: many of the test datasets are of widespread use and the simplest answer is that modern LLMs are trained on this evaluation data. We note that intentional (i.e., to cheat) or unintentional contamination has the same net effect: training on evaluation data limits the confidence and utility of the benchmark.
Curating high-quality datasets of natural language to code instructions can incur exorbitant costs when one example may cost upwards of dozens of US dollars. For any party considering the Pareto optimality of dataset size and coding performance gain, the required funding to create novel data can quickly explode. This leads to a common practice of web scraping code-oriented resources (e.g., GitHub or Stack Overflow) for data. However, these resources are also likely sources of contamination. The small data size and portability of such benchmarks encourages replication. We demonstrate this proliferation by keyword searching for HumanEval prompts on GitHub. Fig. 1 shows that we return a hit in all cases—the median hits is $99$ and the minimum $43$. These hits are often exact duplicates indicating a fork of the original dataset.
Figure 1: Histogram (excluding outliers) of occurrences for HumanEval prompts in public GitHub repositories. Every prompt occurs at least $43$ times.
While decontamination of training sets is becoming more common, present decontamination filters designed for natural text adapt poorly to code. To operate efficiently at scale, most filters rely on generic deduplication algorithms e.g., such as $n$-gram matching or hashing functions Lee et al. (2022). Such surface-level matching does not adequately capture code similarity where a simple variable name change leaves program semantics unchanged, but changing a single keyword can have profound changes. [^1] Compare the instruction “return true if string is float” with “return true if string is verb”. As an example, Elazar et al. (2024) report that only $1.22\%$ of verbatim HumanEval is present in the OSCAR popular web corpus. The same shortcomings of decontamination apply to the creation of large-scale synthetic datasets: for example, the model-generated dataset of Starcoder (Li et al., 2023) is decontaminated only by removing exact docstrings or solutions that match HumanEval or MBPP.
The recent exploration of Riddell et al. (2024) quantifies the proportion of this data leakage in existing datasets using plagiarism tools specifically designed for code. Even when static training datasets are cleaned, contamination may persist through incremental leakage in other sources. For example, entities serving models via an API may encounter these benchmark examples when evaluated by third-party users. When a sample of real model usage is annotated for future training data, samples from benchmark evaluation can leak into future training corpora. Furthermore, these samples may include paraphrases and format changes that further complicate heuristic deduplication. In this scenario, a model may easily memorize completions to purportedly novel prompts. As evidence of this phenomenon, we prompted one popular commercial system (kept anonymous) with partial prompts from HumanEval that were designed to keep the instruction under-specified. Table A in Appendix A shows the outcome and evidence that—despite the ambiguity of the prompt—the result matches exactly the gold solution from HumanEval.
Figure 2: Histogram of cosine similarities for prompts in HumanEval, MBPP and LBPP relative to two popular synthetic code training datasets. We note the high similarity between most HumanEval prompts to evol-instruct, and how LBPP has reduced overall similarity to either training dataset.
3.2 Data leakage through synthetic data
The most capable of code language models rely heavily on the use of synthetic training data (Xu et al., 2023; Wei et al., 2023, 2024; Llama Team, 2024). A typical pipeline consists of curating prompts related to code generation, inferring completions with a previously trained LLM, and synthesizing unit tests for relevant prompts using LLMs. Completions passing the respective unit tests are considered valid code solutions and can be used as future training examples. Alternatively, if a sufficiently powerful model is used, completions might be used as-is without validation.
Synthetic data unlocks scales that are usually not reachable with human-labeled data. Common synthetic code datasets generally have between tens of thousands and millions of examples. For example, Starcoder2-Instruct is a code dataset of around 238k instances (prior to deduplication) that was created by sampling code from GitHub and using it as seeds to generate self-contained code problems, solutions, and associated tests. evol-instruct is another widely popular dataset used by code LLMs such as WizardCoder (Xu et al., 2023) for training. It comprises $110k$ complex query prompts with non-verified completions from closed and open-source models. [^2] Per downloads, the most popular version is a ‘lightly decontaminated’ version on HuggingFace here. The sheer size of these datasets – compared to the domain they are targeting – might explain some memorization. After all, the number of unique, self-contained interview-like prompts with a reasonable size is fairly limited, and it is possible that synthetic datasets cover a majority of this space. When training data covers most examples of a domain, it does not matter whether the model memorizes the training data or whether it can generalize further. Table 3 shows some examples of very similar (but not necessarily equivalent) data between evol-instruct and MBPP.
However, the use of a synthetic data pipeline might hide real leakage. Prior reports  (Yu et al., 2023, page 8)(Wei et al., 2023, page 4) discuss an apparent high similarity between examples in evol-instruct and HumanEval. We also found a lot of semantically equivalent prompts between these two datasets and displayed some in Table A. We extend this analysis by studying the similarity between embedded representations [^3] Embedded using Cohere embed v3 (Team, 2024). of the prompts of HumanEval and MBPP with nearest neighbors from evol-instruct and Starcoder-Instruct. Fig. 2 highlights a widespread similarity between synthetic training datasets and public evaluation data, while the similarity with LBPP is uniformly lower. This is despite LBPP having a very similar format to MBPP and HumanEval (short prompts asking to solve logic problems). The main difference between MBPP/ HumanEval and LBPP is the difficulty level: LBPP’s questions are generally harder (see Section 4). This inference aligns with observations when fine-tuning a ‘Command R Refresh’ model adding evol-instruct. HumanEval score increases by +9%, MBPP increases by +2%, but the LBPP score is unchanged.
Whether the high similarity of synthetic datasets with public evaluation data is due to synthetic data filling the space of problems similar to HumanEval/ MBPP or more direct leakage (eg, through the use of in-context examples), these results point to a larger issue. HumanEval/ MBPP cannot be used as the only proxies to evaluate a model’s code abilities. They mostly provide code performance signal on a very specific type of problems with a very specific level of difficulty. We need more diversity in the code evaluation benchmarks and we believe that LBPP is a step in the right direction.
3.3 Overfitting to test sets
The exaggerated importance of a few benchmarks encourages an incentive structure where model selection prioritizes gain on a narrow suite of metrics. While it may be tempting to use such benchmarks as a deciding factor between similar checkpoints, section 3.2 shows that the correlation between these benchmarks and ‘solving code generation’ is weak. While the meaning and measurement of this unscientific objective are subject to constant revision, selecting for optimal HumanEval performance is akin to $p$-hacking in other fields. This can be justified by assuming these benchmarks are the new dev sets, while the true test is the usage of users over time. However, the usefulness of a dev set entirely relies on its similarity with the actual use case.
Moreover, some risk remains that models overfit to these ‘lucrative’ benchmarks, distorting the perception of downstream performance. Table 2 and Fig. 3 illustrate this problem well. Even though the correlation between MBPP/HumanEval scores and LBPP scores is strong, some models are ranking noticeably higher on MBPP/HumanEval than on LBPP.
The ultimate effect is imbalanced optimization solely towards these metrics, further motivating the practices outlined in Sections 3.1 and 3.2.
Unsolved problems in LBPP
Given a list of unique words each of size $k$
and an $n$ sized word, w, where n is a multiple of $k$,
write a program in Python to determine the number
of unique combinations of words in the list that can
be concatenated to form an anagram of the word.
Write a function in python that takes as input a
recursive function and some parameters. The function
should return the number of times the recursive function
ran itself when starting it with the provided parameters.
Write a Monte Carlo function in Python to compute
the median number of cards you’d need to draw
from a deck such that the sum equals or exceeds
the value $V$.
Table 1: Sampled prompts in LBPP unsolved by leaders on existing benchmarks. Prompts shortened for brevity.
4 LBPP: Less Basic Python Problems
As mentioned above, we have created Less Basic Python Problems (LBPP) for a less biased measure of modern models’ code capabilities. LBPP is a dataset of 161 code completion problems in the style of HumanEval. All prompts include evaluation unit tests, with a median of 4 tests per example.
Dataset Annotation:
Human annotators were asked to create brand new problems that were not solvable by a strong internal model in the loop. All annotators had competitive programming experience. They were instructed to come up with unique problems either from scratch or inspired by programming textbooks whose content was not freely available (e.g., searchable or unlicensed) on the Internet. Annotators could not copy any exercise from a web source or any LLM and only use these sources for inspiration. All sources were cited by annotators and prompts were verified to not match source inspiration. Every prompt is also manually verified as not easily searchable on the web at the time of writing. Each prompt went through additional review to ensure that they were original, hard, and unambiguous. Around a third of the suggested prompts were disqualified for one of these reasons.
All annotators were paid above minimum wage in their respective countries, and all final prompt-completion pairs were manually reviewed by the authors. This adversarial collection resulted in more difficult problems, with most models solving less than $50\%$ of the dataset.
[…]
Mistral Mistral 7B 0.31 0.32 0.11 27$\rightarrow$26
Mixtral $8\times 7$B 0.53 0.23 0.17 22$\rightarrow$23
Mixtral $8\times 22$B 0.73 0.64 0.38 13$\rightarrow$11
Mistral Large 0.92 0.74 0.50 1$\rightarrow$5
Codestral 22B 0.82 0.48 0.40 5$\rightarrow$9
Meta Codellama 7B Instruct 0.39 0.37 0.14 25$\rightarrow$24
Codellama 34B Instruct 0.53 0.53 0.19 23$\rightarrow$22
Llama2 7B Chat 0.17 0.19 0.02 28$\rightarrow$28
Llama2 60B Chat 0.32 0.31 0.10 26$\rightarrow$27
Llama3 8B Instruct 0.62 0.44 0.27 20$\rightarrow$18
Llama3 70B Instruct 0.82 0.67 0.49 6$\rightarrow$6
OpenAI GPT-3.5 Turbo 0.75 0.70 0.37 9$\rightarrow$12
GPT-4 0.82 0.80 0.53 7$\rightarrow$4
GPT-4o 0.90 0.80 0.63 2$\rightarrow$2
Antropic Claude-2 0.65 0.39 0.34 17$\rightarrow$15
Claude-3-Haiku 0.77 0.64 0.34 8$\rightarrow$14
Claude-3-Sonnet 0.74 0.66 0.40 10$\rightarrow$8
Claude-3-Opus 0.84 0.75 0.54 4$\rightarrow$3
Claude-3.5-Sonnet 0.88 0.78 0.64 3$\rightarrow$1
Qwen Qwen1.5 72B Chat 0.63 0.53 0.20 19$\rightarrow$21
Qwen1.5 110B Chat 0.73 0.57 0.30 12$\rightarrow$16
Qwen 2 72B Instruct 0.74 0.67 0.42 11$\rightarrow$7
Cohere Command R 0.43 0.46 0.12 24$\rightarrow$25
Command R (Refresh) 0.71 0.55 0.35 15$\rightarrow$13
Command R+ 0.65 0.61 0.22 18$\rightarrow$20
Command R+ (Refresh) 0.68 0.61 0.29 16$\rightarrow$17
Deepseek Coder 33B Instr. 0.73 0.66 0.39 14$\rightarrow$10
Databricks DBRX Instr. 0.60 0.57 0.25 21$\rightarrow$19
Table 2: Pass@1 results across popular models for zero-shot HumanEval, MBPP and LBPP. All models perform worse on LBPP than either existing benchmark. Rankings between models change between HumanEval and LBPP, contrasting to Riddell et al. (2021). Model rankings similarly also change between MBPP and LBPP.
Initial Results:
Table 2 shows Pass@1 on LBPP for a range of models. Notably, leading models for HumanEval and MBPP perform up to $43\%$ and $27\%$ poorer on LBPP respectively. Additionally, model rankings between either HumanEval and MBPP update for LBPP, potentially identifying overfitting to public benchmarks when presented with a challenging, unseen test set. LBPP is a similarly reliable indicator of code generation performance than prior benchmarks. In Fig. 3, we observe a strong significant correlation between ‘Pass@1’ scores of either HumanEval or MBPP and LBPP. While existing public benchmarks are still a valuable target signal for performance, LBPP is additionally advantageous in that the problems are harder (see Table 2), the dataset is uncontaminated in current training corpora, and prompts bear lower resemblance to existing synthetic corpora (see Fig. 2).
(a) HumanEval
(b) MBPP
Figure 3: Pass@1 rate of LBPP against (a) HumanEval and (b) MBPP. LBPP performance correlates with both prior datasets, but is designed to be genuinely unseen by contemporary LLMs.
Challenges in LBPP:
We study common errors and mistakes from multiple models to identify the most challenging features in LBPP. Examples of problems unsolved by all models are shown in Table 1. Considering common errors between Claude 3.5-Sonnet and Command R Refresh, as the best and recently released model respectively, we identify multiple core trends in failure. Of mutual errors: 21% are related to problems on 2D and 3D arrays; 18% are related to graph-oriented algorithms; and 17% are concerning complex programming concepts often used in competition settings. Additional challenging topics include bit arithmetic & manipulation (8%), Pandas data processing (8%), and file IO (8%). The range of shortcomings between all models highlights the variety of domains that future LLMs must master to improve code generation. Overall, the diversity and difficulty of the problems in LBPP challenges even purportedly advanced models with novel and unsolved prompts.
5 Conclusion
We study the cause and effect of data contamination via two popular code generation benchmarks. Our analysis highlights that contamination is likely unavoidable at the LLM scale given the difficulty of filtering every potential permutation of a benchmark dataset. This insight motivates our contribution of LBPP: a novel code generation benchmark to evaluate contemporary LLMs in a contamination-free setting. We are well aware that our decision to release this dataset will make future leakage impossible to control. However, with the context of the fast-paced model development cycles that LLMs are currently undergoing we believe that releasing this increases the trustworthiness and usefulness of this dataset. It is conveniently designed to serve as a drop-in replacement (or addition) of current evaluation sets. On top of its novelty, the more challenging nature of this dataset also provides a cleaner signal for model comparison.
6 Limitations
All the model analysis was done black-box, without inspecting the model weights or the training set (except the work on synthetic data). There is no reason why this dataset will not follow the same path as the two studied here. As mentioned in the Conclusion we believe there is more value in that than in an alternative solution (not releasing or keeping it behind an API access).
Appendix A Appendix
MBPP test prompt evol-instruct prompt
Write a function to find the
perimeter of a square
Calculate the perimeter of a square
Write a python function to find
the volume of a triangular prism
How can I use JavaScript to
calculate the volume of a
triangular prism with an irregular
base and uneven side lengths?
Write a function to convert snake
case string to camel case string.
Convert the sentence to camel case.
Table 3: Similarity in prompts between MBPP evaluation dataset and evol-instruct synthetic training dataset.
Table 4: Original human evaluation prompts with the completion from a major LLM provider.
Original human-eval prompt Submitted Prompt Opus completion
def is_happy(s): """You are given a string s. Your task is to check if the string is happy or not. A string is happy if its length is at least 3 and every 3 consecutive letters are distinct For example: is_happy(a) => False is_happy(aa) => False is_happy(abcd) => True is_happy(aabb) => False is_happy(adb) => True is_happy(xyy) => False """ Continue this code. Do not add any test, just continue it:
def is_happy(s): """ You are given a string s. Your task is to check if the str
Here’s the continuation of the code without adding any tests: def is_happy(s): """ You are given a string s. Your task is to check if the string is happy. A string is called happy if its length is at least 3 and every 3 consecutive letters are distinct [...]
def count_upper(s): """ Given a string s, count the number of uppercase vowels in even indices. For example: count_upper(’aBCdEf’) returns 1 count_upper(’abcdefg’) returns 0 count_upper(’dBBE’) returns 0 """ Complete this instruction, that is taken from the Human Eval test set: def count_upper(s): """ Given a string s, count the number of Here is the completion of the given function signature and docstring to count the number of uppercase letters in a string: def count_upper(s): """ Given a string s, count the number of uppercase vowels in the string. [...]
def any_int(x, y, z): ’’’ Create a function that takes 3 numbers. Returns true if one of the numbers is equal to the sum of the other two, and all numbers are integers. Returns false in any other cases. Examples: any_int(5, 2, 7) -> True any_int(3, 2, 2) -> False any_int(3, -2, 1) -> True any_int(3.6, -2.2, 2) -> False ’’’ continue this code. def any_int(x, y, z): """Create a function that takes 3 numbers. Return true i def any_int(x, y, z): """ Create a function that takes 3 numbers. Returns true if one of the numbers is equal to the sum of the other two numbers. Otherwise returns false. """ [...]
def order_by_points(nums): """ Write a function which sorts the given list of integers in ascending order according to the sum of their digits. Note: if there are several items with similar sum of their digits, order them based on their index in original list. For example: order_by_points([1, 11, -1, -11, -12]) == [-1, -11, 1, -12, 11] order_by_points([]) == [] """ You will be given an instruction to complete a python function. This instruction is cut in half, so you need to complete it. No need of generating the final code, only the instruction is required. The incomplete instruction is: def order_by_points(nums): """ Write a function which sorts the given list of integers in ascending order according Here’s the completed instruction for the Python function: def order_by_points(nums): """ Write a function which sorts the given list of integers in ascending order according to the sum of their digits. If two integers have the same sum of their digits, order them based on their index in the original list. """
Table 5: Most similar prompt in evol-instruct for a random sample of HumanEval prompts.
HumanEval prompt Closest evol-instrct prompt
Write a Python function longest(strings: List[str]) -> Optional[str]‘ to solve the following problem: Out of list of strings, return the longest one. Return the first one in case of multiple strings of the same length. Return None in case the input list is empty. $>$ longest([]) $>$ longest([’a’, b’, c’]) a $>$ longest([’a’, bb’, ccc’]) ccc Complete the code below, considering an augmented depth of inquiry and maintaining brevity: from typing import List, Optional def longest(strings: List[str]) -> Optional[str]: """ From a list of strings, return the longest one. For multiple strings with equal length, return the first. For an empty list, return None. $>$ longest([]) $>$ longest([’a’, b’, c’]) a $>$ longest([’a’, bb’, ccc’]) ccc
Write a Python function make_a_pile(n)‘ to solve the following problem: Given a positive integer n, you have to make a pile of n levels of stones. The first level has n stones. The number of stones in the next level is: - the next odd number if n is odd. - the next even number if n is even. Return the number of stones in each level in a list, where element at index i represents the number of stones in the level (i+1). Examples: $>$ make_a_pile(3) [3, 5, 7] Please complete the following code with added difficulty: def make_a_pile(n, pattern): """ Given a positive integer n, you have to make a pile of n levels of stones. The first level has n stones. The number of stones in the next level is determined by the given pattern odd or even’: - If pattern is odd’, add the next odd number to the previous level stones. - If pattern is even’, add the next even number to the previous level stones. Return the number of stones in each level in a list, where element at index i represents the number of stones in the level (i+1). Examples: $>$ make_a_pile(3, odd’) [3, 5, 7] $>$ make_a_pile(3, even’) [3, 6, 9] """
Write a Python function x_or_y(n, x, y)‘ to solve the following problem: A simple program which should return the value of x if n is a prime number and should return the value of y otherwise. Examples: for x_or_y(7, 34, 12) == 34 for x_or_y(15, 8, 5) == 5 Complete the subsequent lines of programming code: /* An elementary program intended to, given a single n value, decipher between two distinct possibilities: - If the provided n is a numerical prime, the program should yield the output value equivalent to variable x. - Should the n fail to meet the prime criteria, the program should then present the value contained within variable y as its output. Examples: Upon running x_or_y(7, 34, 12) the output will be 34 While executing x_or_y(15, 8, 5) the output will be 5 */ #include<stdio.h> using namespace std; int x_or_y(int n,int x,int y){